Genuine Spice Inc. began operations on January 1 of the current year. The company produces eight- ounce bottles of hand and bDIRECT LABOR Department Cost Behavior Time per Case Labor Rate per Hour Cost per Case 20 min $18.00 $6.00 Mixing Variable VarPart B-August Budgets During July of the current year, the management of Genuine Spice Inc. asked the controller to prepare AThere was negligible work in process inventory assumed for either the beginning or end of the month; thus, none was assumed.Part C-August Variance Analysis During September of the current year, the controller was asked to perform variance analyses f5.60 min Filling 14.00 Actual variable overhead $305,00 Normal volume 1,600 cases The prices of the materials were differentAmount Descriptions-Part A Controllable variance Equipment depreciation Facility lease Supplies Utilities Volume variance5. Prepare the August production budget. Enter all amounts as positive numbers. Genuine Spice Inc. Production Budget For the6. Prepare the August direct materials purchases budget. Enter all amounts as positive numbers. Genuine Spice Inc. Direct Mat
Answer #1
1 Under high-low method,Highest and lowest level of activity is selected
Highest activity=Case production of 1200 in Mar
Lowest activity=Case production of 500 in Jan
Variable cost per case=Change in total utility cost/Change in cases produced=(740-600)/(1200-500)=140/700=$ 0.20 per case
Fixed cost=Total utility cost-(Cases produced*Variable cost per unit)
Let's take case production of 500
Fixed cost=600-(500*0.20)=600-100=$ 500
2 Contribution margin per case=Selling price-Variable costs
Variable costs:
$
Direct materials 17
Direct labor 7.2
Utilities cost-Variable 0.2
Selling commission 20
Total 44.4
Contribution margin per case=100-44.4=$ 55.6 per case
3 Fixed costs per month:
$
Utilities 500
Facility lease 14000
Equipment depreciation 4300
Supplies 660
Total 19460
4 Break-even number of cases per month=Fixed costs per month/Contribution margin per case=19460/55.6=350 cases
5 Production budget
Cases
Demand for the month 1500
Plus: Desired finished goods inventory 175
Total cases required 1675
Less:Estimated finished goods inventory,August 1 300
Cases to be produced 1375
6 Direct materials purchases budget
Cream base (oz.) Natural Oils (oz.) Bottles (bottles) Total
Cases to be produced a 1375 1375 1375
Material required per case b 100 30 12
Total material required a*b 137500 41250 16500
Plus:Desired materials inventory,August 31 1000 360 240
Less:Estimated materials inventory,August 1 250 290 600
Direct materials to be purchased 138250 41320 16140
x Cost per unit 0.02 0.3 0.5
Direct materials purchase in $ 2765 12396 8070 23231